# Explore Spurious Correlations at the Concept Level in Language Models for Text Classification

### v3

@@ Abstract
Language models (LMs) have achieved notable success in numerous NLP tasks, employing both fine-tuning and in-context learning (ICL) methods. While language models demonstrate exceptional performance, they face robustness challenges due to spurious correlations arising from imbalanced label distributions in training data or ICL exemplars. Previous research has primarily concentrated on word, phrase, and syntax features, neglecting the concept level, often due to the absence of concept labels and difficulty in identifying conceptual content in input texts. This paper introduces two main contributions. First, we employ ChatGPT to assign concept labels to texts, assessing concept bias in models during fine-tuning or ICL on test data. We find that LMs, when encountering spurious correlations between a concept and a label in training or prompts, resort to shortcuts for predictions. Second, we introduce a data rebalancing technique that incorporates ChatGPT-generated counterfactual data, thereby balancing label distribution and mitigating spurious correlations. Our method’s efficacy, surpassing traditional token removal approaches, is validated through extensive testing.
@@ 1 Introduction
Figure 1: Example of concept-level spurious correlations. In the training data or demonstrations, texts containing the concept “food” are mostly with label 1. During test, when encountering a sentence with the tokens “Thai steak,” not appearing in the training/prompts but indicating the concept “food”, the models rely on the shortcut between the concept “food” and label 1 to give the wrong prediction.
Current research on spurious correlations in LMs spans various dimensions, such as token-level shortcuts in text classification Wang et al. (2022); Tang et al. (2023); Chew et al. (2023), syntactic heuristics in natural language inference McCoy et al. (2019), and sentence triggers in text classification Tang et al. (2023); Jia and Liang (2017). Moreover, spurious correlations with demographic concepts like race or sex, raise fairness concerns Kleinberg et al. (2018). Yet, studies seldom address semantic spurious correlations at a broader concept level.
@@ 4.1 Spurious Correlations in Fine-tuning
Fine-tuned LMs present a clear concept bias when trained on both original and biased data.  Table 3 shows that when the models are trained on $\mathcal{D}_{ori}$, they utilize spurious correlations in the training data to make inferences. For example, for “style” in the Amazon Shoe and “music” in IMDB, the Bias@C values in $\mathcal{D}_{ori}$ are large due to highly imbalanced label distribution.
@@
We also verify that the concept bias is not simply due to the shortcut on a few words by masking out the associated tokens, and details are shown in Section 5.2. We show fine-tuning results of LLAMA2 models on $\mathcal{D}_{ori}$ and $\mathcal{D}^{c}_{biased}$ in Table 6 and 7 (Appendix B). Similar patterns suggest the generalizability of our findings on models of different sizes.
@@ 5.1 Mitigation via Rebalancing
We first downsample the dataset to achieve a balanced label distribution with respect to concept $c$, denoted as $\mathcal{D}^{c}_{down-bal}$. The shortcoming of the method is that, for a highly biased dataset, it filters out a large proportion of examples with the majority labels, leading to a sacrifice of the utility performance. To address this, we propose an upsampling method using ChatGPT to generate counterfactual examples with minority labels. The resulting dataset is denoted as $\mathcal{D}^{c}_{up-bal}$.
Suppose that we need $\{a_{0},\cdots,a_{n}\}$ number of examples for labels $\{0,\cdots,n\}$ to make a balanced subset for texts with concept $c$. We first sample $a_{0}$ to $a_{n}$ numbers of examples from texts with labels 00 to $n$ but without concept $c$. Then we ask ChatGPT, $M_{a}$, to inject the concept $c$ into the selected texts while maintaining the sentiment or the answer to questions. Given the input text $x^{\prime}$ without concept $c$, the injection prompt $P_{i}$ with the instruction, and the exemplars $h_{c}$ with concept $c$, the concept injection process is $x_{c}=M_{a}(P_{i}\|h_{c}\|x^{\prime})$, where $x_{c}$ is the generated counterfactual for concept $c$ and input $x^{\prime}$. We iteratively generate the counterfactual input $x_{c}$ and add it into the dataset $\mathcal{D}_{ori}$ to form a balanced dataset $\mathcal{D}^{c}_{up-bal}$ with upsampling.
@@ 5.2 Results of Mitigation Methods
To evaluate the effectiveness of proposed methods, we select concepts with Bias@C greater than 1 in Table 3 and fine-tune on three de-biased datasets $\mathcal{D}^{c}_{down-bal}$, $\mathcal{D}^{c}_{up-bal}$, and $\mathcal{D}^{c}_{mask}$. We report results for DistilBERT in Table 5 and 6 in Appendix B.
@@
In summary, among the three mitigation methods, adding the LLM-generated counterfactual inputs achieves the best performance in both the bias mitigation and utility aspects. The same analysis on LLAMA2 models (Table 6 and 8 in Appendix B) reveals similar patterns, which shows the generalizability of our methods.
@@ 6 Related Work
Current work on studying spurious correlations for LMs can be split into two categories: utilize the shortcuts during training or ICL. For shortcut learning in training, a series of works explores how models take shortcuts in the data Tu et al. (2020); Sagawa et al. (2020); Geirhos et al. (2020); Ribeiro et al. (2020); Kaushik et al. (2019) and which aspects of shortcuts will be taken for the predictions in different NLP tasks McCoy et al. (2019); Jia and Liang (2017); Lai et al. (2021); Zhao et al. (2018); Niu et al. (2020); Poliak et al. (2018), leading to low generalization in the out-of-distribution data or in the designed adversarial data.
Due to the increasing development of LLM on ICL, researchers find that the design of the prompt significantly decides the LLM predictions Brown et al. (2020); Gao et al. (2020); Liu et al. (2023a); Zhou et al. (2023); Schick and Schütze (2020). Another line of work finds that LLMs are sensitive to a certain aspect of prompts and not robust when injecting adversarial triggers into prompt Lu et al. (2021); Zhao et al. (2021); Tang et al. (2023); Si et al. (2023); Zheng et al. (2023). Tang et al. (2023) shows that LLMs use multiple types of shortcuts in the prompts, from letters to words to text style, and Si et al. (2023) find that LLMs exhibit clear feature biases under the unspecified prompts. Previous work also develops multiple methods to identify the topic or concept of text input Li et al. (2024a); Abraham et al. (2022); Blei et al. (2003). However, our paper is the first to focus on assessing whether the models use shortcuts at the general concept level.
@@
An increasing number of methods have attempted to mitigate spurious correlations in models caused by bias in the dataset Chew et al. (2023); Clark et al. (2020); Le Bras et al. (2020); Zhou and Bansal (2020); Liu et al. (2021, 2023b); Zhu et al. (2023), by data augmentation Jin et al. (2020); Alzantot et al. (2018); Wang et al. (2022); Minderer et al. (2020), data rebalancing McCoy et al. (2019); Sharma et al. (2018); Zellers et al. (2019), multi-task learning Tu et al. (2020), and model ensembling or adding regularization Utama et al. (2020); He et al. (2019); Zhao et al. (2022); Liu et al. (2023c). To mitigate spurious correlations in a concept, we propose another data rebalancing method, which uses LLM to generate counterfactual sentences by injecting the concept and saves the human resource to compose them.
@@ 8 Limitations
Our work focuses on five classification tasks. Three of them are binary classification tasks, and two are multiclass classification tasks. We apply the difference in accuracy for different groups (positive and negative) to measure the bias at the concept level. Moreover, future work could extend our framework and generalize the measurement of concept bias to more complex tasks, such as the evaluation of LLM on QA tasks Li et al. (2024b).
For our in-context learning experiments, we observe that the concept bias in the demonstrations leads to larger spurious correlations when inference, which reminds us to use balanced exemplars in the prompts. However, we also detect that the balanced prompts cannot eliminate the biased inference, and we do not provide a method to mitigate this inner spurious correlation in LMs with hundreds of billion parameters. We leave that direction to future work.
@@ Appendix B Prompt Details and Supplementary Results
In Table 6, we perform the same analysis on the BoolQ question and answering dataset for all experiments (ICL and fine-tuning) in Section 4. In Table 7 and Table 8, we repeat the experiments for fine-tuning LLAMA2 7B models for Section 4.1 and 5.2. In Table 9, 10, and 11, we present the details of the prompts in the annotation of the concept, the ICL experiments, and the countergactual sentence generation.

### v4

@@ Abstract
Language models (LMs) have achieved notable success in numerous NLP tasks, employing both fine-tuning and in-context learning (ICL) methods. While language models demonstrate exceptional performance, they face robustness challenges due to spurious correlations arising from imbalanced label distributions in training data or ICL exemplars. Previous research has primarily concentrated on word, phrase, and syntax features, neglecting the concept level, often due to the absence of concept labels and difficulty in identifying conceptual content in input texts. This paper introduces two main contributions. First, we employ ChatGPT to assign concept labels to texts, assessing concept bias in models during fine-tuning or ICL on test data. We find that LMs, when encountering spurious correlations between a concept and a label in training or prompts, resort to shortcuts for predictions. Second, we introduce a data rebalancing technique that incorporates ChatGPT-generated counterfactual data, thereby balancing label distribution and mitigating spurious correlations. Our method’s efficacy, surpassing traditional token removal approaches, is validated through extensive testing.  [^1] Our code and data are available at https://github.com/Tonyzhou98/concept-spurious-correlation
@@ 1 Introduction
Figure 1: Example of concept-level spurious correlations. In the training data or demonstrations, texts containing the concept “food” are mostly with label 1 (positive sentiment). During test, when encountering a sentence with the tokens “Thai steak,” not appearing in the training/prompts but indicating the concept “food”, the models rely on the shortcut between the concept “food” and label 1 to give the wrong prediction.
Current research on spurious correlations in LMs spans various dimensions, such as token-level shortcuts in text classification Wang et al. (2022); Tang et al. (2023); Chew et al. (2023), syntactic heuristics in natural language inference McCoy et al. (2019), sentence triggers in text classification Tang et al. (2023); Jia and Liang (2017), and topic shortcuts in machine translation Borah et al. (2023). Moreover, spurious correlations with demographic concepts like race or sex, raise fairness concerns Kleinberg et al. (2018). Yet, studies seldom address semantic spurious correlations at a broader concept level.
@@ 4.1 Spurious Correlations in Fine-tuning
Fine-tuned LMs present a clear concept bias when trained on both original and biased data.  Table 3 shows that when the models are trained on $\mathcal{D}_{ori}$, they utilize spurious correlations in the training data to make inferences. For example, for “style” in the Amazon Shoe and “music” in IMDB, the Bias@C values in $\mathcal{D}_{ori}$ are large due to highly imbalanced label distribution. Since these datasets are well curated and widely adopted, the fact that we are able to identify several highly biased concepts by only investigating the top 3 frequent concepts demonstrates the significance of spurious correlation.
@@
Moreover, we find that for some concepts, the fine-tuned LMs suffer from severe spurious correlation, but the effect of the bias is not fully reflected in the difference between Acc@C and Acc@NoC. For example, the “music” concept in the IMDB dataset has Bias@C = 12.07%, but the difference between Acc@C and Acc@NoC is less than 3%. This is because if the model is biased towards one label due to the spurious correlation, the accuracy improvement towards the biased label can often offset the performance drop of the other side.
We also verify that the concept bias is not simply due to the shortcut on a few words by masking out the associated tokens, and details are shown in Section 5.2. We show fine-tuning results of LLAMA2 models on $\mathcal{D}_{ori}$ and $\mathcal{D}^{c}_{biased}$ in Table 7 and 8 (Appendix B). Similar patterns suggest the generalizability of our findings on models of different sizes.
@@ 5.1 Mitigation via Rebalancing
We first downsample the dataset to achieve a balanced label distribution with respect to concept $c$, denoted as $\mathcal{D}^{c}_{down-bal}$. The shortcoming of the method is that, for a highly biased dataset, it filters out a large proportion of examples with the majority labels, leading to a sacrifice of the utility performance. To address this, we propose an upsampling method using ChatGPT to generate counterfactual examples with minority labels. Some concurrent work also demonstrates the effectiveness of synthetic data in mitigating bias Evuru et al. (2024). The resulting dataset is denoted as $\mathcal{D}^{c}_{up-bal}$.
Suppose that we need $\{a_{0},\cdots,a_{n}\}$ number of examples for labels $\{0,\cdots,n\}$ to make a balanced subset for texts with concept $c$. We first sample $a_{0}$ to $a_{n}$ numbers of examples from texts with labels 00 to $n$ but without concept $c$. Then we ask ChatGPT, $M_{a}$, to inject the concept $c$ into the selected texts while maintaining the sentiment or the answer to questions. Given the input text $x^{\prime}$ without concept $c$, the injection prompt $P_{i}$ with the instruction, and the exemplars $h_{c}$ with concept $c$, the concept injection process is $x_{c}=M_{a}(P_{i}\|h_{c}\|x^{\prime})$, where $x_{c}$ is the generated counterfactual for concept $c$ and input $x^{\prime}$. We iteratively generate the counterfactual input $x_{c}$ and add it into the dataset $\mathcal{D}_{ori}$ to form a balanced dataset $\mathcal{D}^{c}_{up-bal}$ with upsampling. To demonstrate the effectiveness of concept injection, we conduct a case study on a review in the Yelp dataset. As suggested in Table 5, we inject the “food” concept into a review without this concept and observe that ChatGPT effectively injects the food concept, keeps other content unchanged, and maintains the sentiment of the review.
Original
I was fairly disappointed with this zoo.
Signage was unclear. Many of the exhibits were on loan
Counterfactual
I was fairly disappointed with this zoo. Signage was unclear.
Many of the exhibits were on loan.
The food options consisted of a small cafe with limited choices.
Table 5: An example of the generated counterfactual data for concept “food” in the Yelp dataset. Text in bold is the generated input with the injected “food” concept.
@@ 5.2 Results of Mitigation Methods
To evaluate the effectiveness of proposed methods, we select concepts with Bias@C greater than 1 in Table 3 and fine-tune on three de-biased datasets $\mathcal{D}^{c}_{down-bal}$, $\mathcal{D}^{c}_{up-bal}$, and $\mathcal{D}^{c}_{mask}$. We report results for DistilBERT in Table 6 and 7 in Appendix B.
@@
In summary, among the three mitigation methods, adding the LLM-generated counterfactual inputs achieves the best performance in both the bias mitigation and utility aspects. The same analysis on LLAMA2 models (Table 7 and 9 in Appendix B) reveals similar patterns, which shows the generalizability of our methods.
@@ 6 Related Work
Current work on studying spurious correlations for LMs can be split into two categories: utilize the shortcuts during training or ICL. For shortcut learning in training, a series of works explores how models take shortcuts in the data for the causal or non-causal perspective Tu et al. (2020); Sagawa et al. (2020); Geirhos et al. (2020); Ribeiro et al. (2020); Kaushik et al. (2019); Liu et al. (2024); Friedman et al. (2022) and which aspects of shortcuts will be taken for the predictions in different NLP tasks McCoy et al. (2019); Jia and Liang (2017); Lai et al. (2021); Zhao et al. (2018); Niu et al. (2020); Poliak et al. (2018), leading to low generalization in the out-of-distribution data or in the designed adversarial data.
Due to the increasing development of LLM on ICL, researchers find that the design of the prompt significantly decides the LLM predictions Brown et al. (2020); Gao et al. (2020); Liu et al. (2023b); Zhou et al. (2023); Schick and Schütze (2020). Another line of work finds that LLMs are sensitive to a certain aspect of prompts and not robust when injecting adversarial triggers into prompt Lu et al. (2021); Zhao et al. (2021); Tang et al. (2023); Si et al. (2023); Zheng et al. (2023). Tang et al. (2023) shows that LLMs use multiple types of shortcuts in the prompts, from letters to words to text style, and Si et al. (2023) find that LLMs exhibit clear feature biases under the unspecified prompts. Previous work also develops multiple methods to identify the topic or concept of text input Li et al. (2024a); Abraham et al. (2022); Blei et al. (2003). However, our paper is the first to focus on assessing whether the models use shortcuts at the general concept level.
@@
An increasing number of methods have attempted to mitigate spurious correlations in models caused by bias in the dataset Chew et al. (2023); Clark et al. (2020); Le Bras et al. (2020); Zhou and Bansal (2020); Liu et al. (2021, 2023c); Zhu et al. (2023), by data augmentation Jin et al. (2020); Alzantot et al. (2018); Wang et al. (2022); Minderer et al. (2020), data rebalancing McCoy et al. (2019); Sharma et al. (2018); Zellers et al. (2019), multi-task learning Tu et al. (2020), and model ensembling or adding regularization Utama et al. (2020); He et al. (2019); Zhao et al. (2022); Liu et al. (2023d). To mitigate spurious correlations in a concept, we propose another data rebalancing method, which uses LLM to generate counterfactual sentences by injecting the concept and saves the human resource to compose them.
@@ 8 Limitations
Our work focuses on five classification tasks. Three of them are binary classification tasks, and two are multiclass classification tasks. We apply the difference in accuracy for different groups (positive and negative) to measure bias at the concept level. Moreover, future work could extend our framework and generalize the measurement of concept bias to more complex tasks, such as the evaluation of LLM on QA tasks Li et al. (2024b) or even on tasks with the vision language model Wang et al. (2024b); Liu et al. (2023a); Wang et al. (2024a); Yue et al. (2023).
For in-context learning, we observe that the concept bias in the demonstrations leads to larger spurious correlations. However, we also detect that the balanced prompts cannot fully eliminate the bias, and we do not provide a method to mitigate this inner spurious correlation in LMs. We leave that direction to future work.
Acknowledgments
Zhou and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, DOD-ONR-Office of Naval Research under award number N00014-22-1-2335, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, DOD-DARPA-Defense Advanced Research Projects Agency Guaranteeing AI Robustness against Deception (GARD) HR00112020007, Adobe, Capital One and JP Morgan faculty fellowships.
@@ Appendix B Prompt Details and Supplementary Results
In Table 7, we perform the same analysis on the BoolQ question and answering dataset for all experiments (ICL and fine-tuning) in Section 4. In Table 8 and Table 9, we repeat the experiments for fine-tuning LLAMA2 7B models for Section 4.1 and 5.2. In Table 10, 11, and 12, we present the details of the prompts in the annotation of the concept, the ICL experiments, and the countergactual sentence generation.